# Optimal transport for types and convex analysis for definable predicates in tracial $\mathrm{W}^{*}$-algebras

### 02 Jan 2024 v2

@@ 1. Introduction
Free probability theory treats tracial (finite) von Neumann algebras as non-commutative probability spaces and investigates the way that ideas from classical probability can be applied in the non-commutative setting. In particular, non-commutative distributions or laws often play the role of probability distributions for a tuple of random variables. The non-commutative analog of Wasserstein distance was introduced in [10], and optimal transport theory has been in the background of several significant developments in free probability and random matrix theory [35, 31, 18, 41]. Non-commutative Monge-Kantorovich duality was addressed in [24]. In [40, §6.1] the present author suggested investigating non-commutative optimal transport theory in the framework of continuous model theory. Continuous model theory [8, 6] made its first formal contact with tracial von Neumann algebras in [30, 21, 22, 23] and has since had many applications [26, 27, 25, 1]; for a survey, see [29]. The non-commutative law of a tuple $(x_{1},\dots,x_{n})$ corresponds to the quantifier-free type in in the model theory of tracial von Neumann algebras described in [22]. However, because diffuse classical probability spaces admit quantifier elimination, there is no distinction between the quantifier-free type and the complete type; see [7, Fact 2.10], [9, §6], [40, §2.3]. Hence, in the non-commutative setting the complete type of a tuple $(x_{1},\dots,x_{n})$ could also be viewed as an analog of the probability distribution. One way that the complete type makes a better analog than the quantifier-free type is the relationship between (complete) types and automorphisms. In the setting of classical probability (or equivalently commutative von Neumann algebras), two tuples $(X_{1},\dots,X_{n})$ and $(Y_{1},\dots,Y_{n})$ of random variables on a diffuse probability space $(\Omega,P)$ have the same probability distribution if and only if they are approximately conjugate by automorphisms of $(\Omega,P)$; see [9, §6], [40, §2.3]. Analogously, every tracial von Neumann algebra $\mathcal{M}$ admits an elementary extension $\mathcal{N}\succeq\mathcal{M}$, such that two tuples $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$ are conjugate by an automorphism of $\mathcal{N}$ if and only if they have the same type (such an $\mathcal{M}$ is called strongly $\aleph_{1}$-homogeneous); see [6, §7]. However, the quantifier-free type does not provide enough information about $(x_{1},\dots,x_{n})$ to determine it up to automorphism in this sense. The analog of the Wasserstein distance for types is known in continuous model theory as the $d$-metric [6, p. 44]. The $L^{1}$-Wasserstein distance in classical probability was studied from a model-theoretic viewpoint by Song [51, §5.3-5.4], while here we handle the $L^{2}$-Wasserstein distance in the non-commutative setting as in [40, §6.1]. In this discussion we fix a complete theory $\mathrm{T}$, i.e. we restrict our attention to one elementary equivalence class of tracial von Neumann algebras. For tuples $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathrm{y}=(y_{1},\dots,y_{n})$, we write $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{j=1}^{n}d^{\mathcal{M}}(x_{j},y_{j})^{2}}=\lVert\mathbf{x}-\mathbf{y}\rVert_{L^{2}(\mathcal{M})^{n}}$. For two types $\mu$ and $\nu$ relative to $\mathrm{T}$, the distance $d_{W}(\mu,\nu)$ is defined as the infimum of $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})$ over all models $\mathcal{M}$ of $\mathrm{T}$ and over all $\mathbf{x},\mathbf{y}\in\mathcal{M}^{n}$ such that $\operatorname{tp}^{\mathcal{M}}(\mathbf{x})=\mu$ and $\operatorname{tp}^{\mathcal{M}}(\mathbf{y})$. We call $(\mathbf{x},\mathbf{y})$ an optimal coupling of $(\mu,\nu)$ if $\operatorname{tp}^{\mathcal{M}}(\mathbf{x})=\mu$ and $\operatorname{tp}^{\mathcal{M}}(\mathbf{y})$ and $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})$ realizes the infimum. If $\mathcal{M}$ is is a $\aleph_{1}$-saturated model of $\mathrm{T}$ (see §2.6), then an optimal coupling of $\mu$ and $\nu$ must exist in $\mathcal{M}$. Moreover, if $\mathcal{M}$ is $\aleph_{1}$-saturated and strongly $\aleph_{1}$-homogeneous (see §2.6), then the types $\mu$ and $\nu$ correspond to $\operatorname{Aut}(\mathcal{M})$-orbits in $\mathcal{M}^{n}$ and $d_{W}(\mu,\nu)$ is the distance between these two orbits. Our first main result is a dual characterization of the Wasserstein distance using definable predicates. In the setting of tracial von Neumann algebras, definable predicates are scalar-valued functions $\phi^{\mathcal{M}}(x_{1},\dots,x_{n})$ that are uniform limits (on each ball) of formulas built out of traces of non-commutative polynomials using continuous functions and $\sup$ and $\inf$ operations. The definable predicates and types are related by a Gelfand duality loosely analogous to $C_{0}(\mathbb{C}^{n},\mathbb{R})$ being the predual of $\mathcal{P}(\mathbb{C}^{n})$. [^1] Technically, in the language of classical probability spaces, the space of definable predicates would be $C(\mathcal{P}(\mathbb{C}^{n}),\mathbb{R})$ rather than $C_{0}(\mathbb{C}^{n},\mathbb{R})$ or even $C(\mathbb{C}^{n},\mathbb{R})$. In the general setting, we have to make do without a direct analog of “points.” Indeed, definable predicates can be identified with continuous functions on the space of types, while conversely, types can be identified with continuous real-valued characters on the space of definable predicates. In classical probability theory, the Wasserstein distance has a dual characterization in terms of continuous functions as follows. Let $\mu$ and $\nu$ be probability distributions on $\mathbb{C}^{n}$ (assume they are compactly supported for simplicity). Let
@@
The case of complete types here works out somewhat better than the case of quantifier-free types in [24]. The $E$-convex functions described there were not necessarily definable predicates, and certainly not always quantifier-free definable predicates. Moreover, we did not necessarily expect to be able to achieve the duality with quantifier-free definable predicates because the weak-*** topology on the space of quantifier-free types is weaker than the metric topology produced by the Wasserstein distance [24, Proposition 1.7]. When using complete types, the analog of Monge-Kantorovich duality works using convex definable predicates, which are continuous with respect to the weak-*** (logic) topology, despite the fact that even for complete types the logic topology is weaker than metric topology from the Wasserstein distance [6, Proposition 12.10]. For $E$-convex functions, one had to assume monotonicity under condition expectations, i.e. $\phi^{\mathcal{N}}(E_{\mathcal{N}}(\mathbf{x}))\leq\phi^{\mathcal{M}}(\mathbf{x})$ for $\mathcal{N}\subseteq\mathcal{M}$ and $\mathbf{x}\in L^{2}(\mathcal{M})^{n}$, as an additional condition beside convexity of $\phi^{\mathcal{M}}$ for each $\mathcal{M}$ [24, Definitions 1.2 and 3.6]. However, for convex definable predicates, monotonicity under conditional expectations onto elementary submodels is automatic.
@@
In [24, Theorem 1.5, §4.3], Monge-Kantorovich duality was used to show that if $\mathbf{x}$ and $\mathbf{y}$ are an optimal coupling of quantifier-free types, then $\mathrm{W}^{*}((1-t)\mathbf{x}+t\mathbf{y})=\mathrm{W}^{*}(\mathbf{x},\mathbf{y})$ for $t\in(0,1)$. In the analogous result for complete types, we replace the von Neumann algebra $\mathrm{W}^{*}(\mathbf{x})$ by the definable closure $\operatorname{dcl}^{\mathcal{M}}(\mathbf{x})$ (see §3 for background).
@@
The definable closure has the following characterization in terms of automorphisms (see Proposition 3.4 for details). Suppose that $\mathcal{M}$ is a strongly $\aleph_{1}$-homogeneous model of $\mathrm{T}$. Then $y$ is in the definable closure of a finite or countable tuple $\mathbf{x}$ if and only if $y$ is fixed by every automorphism of $\mathcal{M}$ that fixes $\mathbf{x}$. Very little is known about the definable closure for von Neumann algebras in general, and thus in §3.2, we give several examples. In particular, we compute the definable closure for all ***-subalgebras of finite-dimensional tracial $\mathrm{W}^{*}$-algebras (Example 3.10). We also show that if $\mathcal{A}$ is a $\mathrm{W}^{*}$-subalgebra of $\mathcal{M}$ with trivial relative commutant and weak spectral gap, then $\operatorname{dcl}^{\mathcal{M}}(\mathcal{A})$ contains the commutator subgroup of the normalizer of $\mathcal{A}$ in $\mathcal{M}$ (Proposition 3.12). In particular, the definable closure is in some cases much larger than the original subalgebra. Our next main result relates the definable closure with definable functions (see Definition 2.19). This result is analogous to [37, Proposition 13.6.6], [34, Proposition 2.4], [39, Propsosition 3.32] which showed that if $A$ is a subset of a tracial von Neumann algebra $\mathcal{M}$, then any $x\in\mathrm{W}^{*}(A)$ can be expressed as a quantifier-free definable function over $\mathrm{T}_{\operatorname{tr}}$ of some $(a_{j})_{j\in\mathbb{N}}$ in $A$.
@@
The challenge of this result is the assertion that there is a continuous selection of an $f((y_{j})_{j\in\mathbb{N}})$ for all $(y_{j})_{j\in\mathbb{N}}$, which when evaluated at $(a_{j})_{j\in\mathbb{N}}$ yields the given $z\in\operatorname{dcl}^{\mathcal{M}}(A)$. The proof is another application of the theory of convex definable predicates. We first observe that if $z$ is in the definable closure, then $\operatorname{Re}\langle z,x\rangle_{L^{2}(\mathcal{M})}=\phi(x,(a_{j})_{j\in\mathbb{N}})$ for some definable predicate $\phi$. We show that any definable predicate $\phi(x,(y_{j})_{j\in\mathbb{N}})$ can be approximated by one which is differentiable with respect to $x$, whose gradient is a definable function; here the gradient is understood by viewing $L^{2}(\mathcal{M})$ as the real Hilbert space with $\operatorname{Re}\langle\cdot,\cdot\rangle$ as the inner product. The gradient is relevant here since $\nabla_{x}\phi(x,(a_{j})_{j\in\mathbb{N}})=\nabla_{x}\operatorname{Re}\langle x,z\rangle=z$; see §5.3 for the proof. Hence, the key ingredient is a regularization technique that allows us to approximate an arbitrary definable predicate by one which is “smooth.” In §5.2, we employ the technique of semiconvex semiconcave regularization by sup-convolutions and inf-convolutions as in [47], which has played a major role in the theory of Hamilton-Jacobi equations on infinite-dimensional Hilbert spaces [2, 4, 3, 15, 16, 17], to achieve the following result.
@@ 2.1. Background on operator algebras
    A $\mathrm{W}^{*}$-algebra is a $\mathrm{C}^{*}$-algebra $\mathcal{A}$ that $\mathcal{A}$ is isomorphic as a Banach space to the dual of some other Banach space (called its predual)
@@ 2.2. Languages and structures
Here let us sketch the setup of continuous model theory, or model theory for metric structures from Ben Yaacov, Berenstein, Henson, and Usvyatsov [6, 8]; see also [32]. We will follow the treatment in Farah, Hart, and Sherman [22] which introduces “domains of quantification” to cut down on the number of “sorts” needed; see also [39, §2-3]. A metric signature $\mathcal{L}$ (sometimes called a language) consists of:
@@
  • For technical reasons (see [22, Proposition 3.2] and [32, p. 93]), we also introduce for each $d$-variable non-commutative polynomial $p$ and $n\in\mathbb{N}$ a symbol $t_{p}:L^{\infty}(\mathcal{M})^{d}$ representing the evaluation of $p$, along with the appropriate range bounds $N_{t_{p},\mathbf{r}}$ between $M_{p}$ and $M_{p}+1/n$, where $M_{p}$ is the supremum of $\lVert p(X_{1},\dots,X_{d})\rVert$ over all $(X_{1},\dots,X_{d})$ in a tracial $\mathrm{W}^{*}$-algebra $\mathcal{M}$.
Each function and relation symbol is assigned range bounds and moduli of continuity that one would expect, e.g. multiplication is supposed to map $D_{r}\times D_{r^{\prime}}$ into $D_{rr^{\prime}}$ with $\omega_{(D_{r},D_{r}^{\prime}),1}^{\cdot}(t)=r^{\prime}t$ and $\omega_{(D_{r},D_{r}^{\prime}),2}^{\cdot}=rt$. Although not every $\mathcal{L}_{\operatorname{tr}}$-structure comes from a tracial $\mathrm{W}^{*}$-algebra, one can formulate axioms in the language such that any structure satisfying these axioms comes from a tracial $\mathrm{W}^{*}$-algebra [22, §3.2]. For such structures, the notion of homomorphism given above reduces to trace-preserving ***-homomorphisms of tracial $\mathrm{W}^{*}$-algebras.

### 09 Mar 2024 v3

@@ 1. Introduction
Free probability theory treats tracial (finite) von Neumann algebras as non-commutative probability spaces and investigates the way that ideas from classical probability can be applied in the non-commutative setting. In particular, non-commutative distributions or laws often play the role of probability distributions for a tuple of random variables. The non-commutative analog of Wasserstein distance was introduced in [10], and optimal transport theory has been in the background of several significant developments in free probability and random matrix theory [36, 32, 18, 42]. Non-commutative Monge-Kantorovich duality was addressed in [25]. In [41, §6.1] the present author suggested investigating non-commutative optimal transport theory in the framework of continuous model theory. Continuous model theory [8, 6] made its first formal contact with tracial von Neumann algebras in [31, 22, 23, 24] and has since had many applications [27, 28, 26, 1]; for a survey, see [30]. The non-commutative law of a tuple $(x_{1},\dots,x_{n})$ corresponds to the quantifier-free type in in the model theory of tracial von Neumann algebras described in [23]. However, because diffuse classical probability spaces admit quantifier elimination, there is no distinction between the quantifier-free type and the complete type; see [7, Fact 2.10], [9, §6], [41, §2.3]. Hence, in the non-commutative setting, many tracial von Neumann algebras do not admit quantifier elimination [21], and so the complete type of a tuple $(x_{1},\dots,x_{n})$ could also be viewed as an analog of the probability distribution. One way that the complete type makes a better analog than the quantifier-free type is the relationship between (complete) types and automorphisms. In the setting of classical probability (or equivalently commutative von Neumann algebras), two tuples $(X_{1},\dots,X_{n})$ and $(Y_{1},\dots,Y_{n})$ of random variables on a diffuse probability space $(\Omega,P)$ have the same probability distribution if and only if they are approximately conjugate by automorphisms of $(\Omega,P)$; see [9, §6], [41, §2.3]. Analogously, every tracial von Neumann algebra $\mathcal{M}$ admits an elementary extension $\mathcal{N}\succeq\mathcal{M}$, such that two tuples $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$ are conjugate by an automorphism of $\mathcal{N}$ if and only if they have the same type (such an $\mathcal{N}$ is called strongly $\aleph_{1}$-homogeneous); see [6, §7]. However, the quantifier-free type does not provide enough information about $(x_{1},\dots,x_{n})$ to determine it up to automorphism in this sense. The analog of the Wasserstein distance for types is known in continuous model theory as the $d$-metric [6, p. 44]. The $L^{1}$-Wasserstein distance in classical probability was studied from a model-theoretic viewpoint by Song [52, §5.3-5.4], while here we handle the $L^{2}$-Wasserstein distance in the non-commutative setting as in [41, §6.1]. In this discussion we fix a complete theory $\mathrm{T}$, i.e. we restrict our attention to one elementary equivalence class of tracial von Neumann algebras. For tuples $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathrm{y}=(y_{1},\dots,y_{n})$, we write $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{j=1}^{n}d^{\mathcal{M}}(x_{j},y_{j})^{2}}=\lVert\mathbf{x}-\mathbf{y}\rVert_{L^{2}(\mathcal{M})^{n}}$. For two types $\mu$ and $\nu$ relative to $\mathrm{T}$, the distance $d_{W}(\mu,\nu)$ is defined as the infimum of $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})$ over all models $\mathcal{M}$ of $\mathrm{T}$ and over all $\mathbf{x},\mathbf{y}\in\mathcal{M}^{n}$ such that $\operatorname{tp}^{\mathcal{M}}(\mathbf{x})=\mu$ and $\operatorname{tp}^{\mathcal{M}}(\mathbf{y})$. We call $(\mathbf{x},\mathbf{y})$ an optimal coupling of $(\mu,\nu)$ if $\operatorname{tp}^{\mathcal{M}}(\mathbf{x})=\mu$ and $\operatorname{tp}^{\mathcal{M}}(\mathbf{y})$ and $d^{\mathcal{M}}(\mathbf{x},\mathbf{y})$ realizes the infimum. If $\mathcal{M}$ is is a $\aleph_{1}$-saturated model of $\mathrm{T}$ (see §2.6), then an optimal coupling of $\mu$ and $\nu$ must exist in $\mathcal{M}$. Moreover, if $\mathcal{M}$ is $\aleph_{1}$-saturated and strongly $\aleph_{1}$-homogeneous (see §2.6), then the types $\mu$ and $\nu$ correspond to $\operatorname{Aut}(\mathcal{M})$-orbits in $\mathcal{M}^{n}$ and $d_{W}(\mu,\nu)$ is the distance between these two orbits. Our first main result is a dual characterization of the Wasserstein distance using definable predicates. In the setting of tracial von Neumann algebras, definable predicates are scalar-valued functions $\phi^{\mathcal{M}}(x_{1},\dots,x_{n})$ that are uniform limits (on each ball) of formulas built out of traces of non-commutative polynomials using continuous functions and $\sup$ and $\inf$ operations. The definable predicates and types are related by a Gelfand duality loosely analogous to $C_{0}(\mathbb{C}^{n},\mathbb{R})$ being the predual of $\mathcal{P}(\mathbb{C}^{n})$. [^1] Technically, in the language of classical probability spaces, the space of definable predicates would be $C(\mathcal{P}(\mathbb{C}^{n}),\mathbb{R})$ rather than $C_{0}(\mathbb{C}^{n},\mathbb{R})$ or even $C(\mathbb{C}^{n},\mathbb{R})$. In the general setting, we have to make do without a direct analog of “points.” Indeed, definable predicates can be identified with continuous functions on the space of types, while conversely, types can be identified with continuous real-valued characters on the space of definable predicates. In classical probability theory, the Wasserstein distance has a dual characterization in terms of continuous functions as follows. Let $\mu$ and $\nu$ be probability distributions on $\mathbb{C}^{n}$ (assume they are compactly supported for simplicity). Let
@@
The case of complete types here works out somewhat better than the case of quantifier-free types in [25]. The $E$-convex functions described there were not necessarily definable predicates, and certainly not always quantifier-free definable predicates. Moreover, we did not necessarily expect to be able to achieve the duality with quantifier-free definable predicates because the weak-*** topology on the space of quantifier-free types is weaker than the metric topology produced by the Wasserstein distance [25, Proposition 1.7]. When using complete types, the analog of Monge-Kantorovich duality works using convex definable predicates, which are continuous with respect to the weak-*** (logic) topology, despite the fact that even for complete types the logic topology is weaker than metric topology from the Wasserstein distance [6, Proposition 12.10]. For $E$-convex functions, one had to assume monotonicity under condition expectations, i.e. $\phi^{\mathcal{N}}(E_{\mathcal{N}}(\mathbf{x}))\leq\phi^{\mathcal{M}}(\mathbf{x})$ for $\mathcal{N}\subseteq\mathcal{M}$ and $\mathbf{x}\in L^{2}(\mathcal{M})^{n}$, as an additional condition beside convexity of $\phi^{\mathcal{M}}$ for each $\mathcal{M}$ [25, Definitions 1.2 and 3.6]. However, for convex definable predicates, monotonicity under conditional expectations onto elementary substructures is automatic.
@@
In [25, Theorem 1.5, §4.3], Monge-Kantorovich duality was used to show that if $\mathbf{x}$ and $\mathbf{y}$ are an optimal coupling of quantifier-free types, then $\mathrm{W}^{*}((1-t)\mathbf{x}+t\mathbf{y})=\mathrm{W}^{*}(\mathbf{x},\mathbf{y})$ for $t\in(0,1)$. In the analogous result for complete types, we replace the von Neumann algebra $\mathrm{W}^{*}(\mathbf{x})$ by the definable closure $\operatorname{dcl}^{\mathcal{M}}(\mathbf{x})$ (see §3 for background).
@@
The definable closure has the following characterization in terms of automorphisms (see Proposition 3.4 for details). Suppose that $\mathcal{M}$ is a strongly $\aleph_{1}$-homogeneous model of $\mathrm{T}$. Then $y$ is in the definable closure of a finite or countable tuple $\mathbf{x}$ if and only if $y$ is fixed by every automorphism of $\mathcal{M}$ that fixes $\mathbf{x}$. Very little is known about the definable closure for von Neumann algebras in general, and thus in §3.2, we give several examples. In particular, we compute the definable closure for all ***-subalgebras of finite-dimensional tracial $\mathrm{W}^{*}$-algebras (Example 3.10). We also show that if $\mathcal{A}$ is a $\mathrm{W}^{*}$-subalgebra of $\mathcal{M}$ with trivial relative commutant and weak spectral gap, then $\operatorname{dcl}^{\mathcal{M}}(\mathcal{A})$ contains the commutator subgroup of the normalizer of $\mathcal{A}$ in $\mathcal{M}$ (Proposition 3.12). In particular, the definable closure is in some cases much larger than the original subalgebra. Our next main result relates the definable closure with definable functions (see Definition 2.19). This result is analogous to [38, Proposition 13.6.6], [35, Proposition 2.4], [40, Propsosition 3.32] which showed that if $A$ is a subset of a tracial von Neumann algebra $\mathcal{M}$, then any $x\in\mathrm{W}^{*}(A)$ can be expressed as a quantifier-free definable function over $\mathrm{T}_{\operatorname{tr}}$ of some $(a_{j})_{j\in\mathbb{N}}$ in $A$.
@@
The challenge of this result is the assertion that there is a continuous selection of an $f((y_{j})_{j\in\mathbb{N}})$ for all $(y_{j})_{j\in\mathbb{N}}$, which when evaluated at $(a_{j})_{j\in\mathbb{N}}$ yields the given $z\in\operatorname{dcl}^{\mathcal{M}}(A)$. The proof is another application of the theory of convex definable predicates. We first observe that if $z$ is in the definable closure, then $\operatorname{Re}\langle z,x\rangle_{L^{2}(\mathcal{M})}=\phi(x,(a_{j})_{j\in\mathbb{N}})$ for some definable predicate $\phi$. We show that any definable predicate $\phi(x,(y_{j})_{j\in\mathbb{N}})$ can be approximated by one which is differentiable with respect to $x$, whose gradient is a definable function; here the gradient is understood by viewing $L^{2}(\mathcal{M})$ as the real Hilbert space with $\operatorname{Re}\langle\cdot,\cdot\rangle$ as the inner product. The gradient is relevant here since $\nabla_{x}\phi(x,(a_{j})_{j\in\mathbb{N}})=\nabla_{x}\operatorname{Re}\langle x,z\rangle=z$; see §5.3 for the proof. Hence, the key ingredient is a regularization technique that allows us to approximate an arbitrary definable predicate by one which is “continuously differentiable.” In §5.2, we employ the technique of semiconvex semiconcave regularization by sup-convolutions and inf-convolutions as in [48], which has played a major role in the theory of Hamilton-Jacobi equations on infinite-dimensional Hilbert spaces [2, 4, 3, 15, 16, 17], to achieve the following result.
@@ 2.1. Background on operator algebras
    A $\mathrm{W}^{*}$-algebra is a $\mathrm{C}^{*}$-algebra $\mathcal{A}$ that $\mathcal{A}$ is isomorphic as a Banach space to the dual of some other Banach space (called its predual).
@@ 2.2. Languages and structures
Here let us sketch the setup of continuous model theory, or model theory for metric structures from Ben Yaacov, Berenstein, Henson, and Usvyatsov [6, 8]; see also [33]. We will follow the treatment in Farah, Hart, and Sherman [23] which introduces “domains of quantification” to cut down on the number of “sorts” needed; see also [40, §2-3]. A metric signature $\mathcal{L}$ (sometimes called a language) consists of:
@@
  • For technical reasons (see [23, Proposition 3.2] and [33, p. 93]), we also introduce for each $d$-variable non-commutative polynomial $p$ and $n\in\mathbb{N}$ a symbol $t_{p}:L^{\infty}(\mathcal{M})^{d}$ representing the evaluation of $p$, along with the appropriate range bounds $N_{t_{p},\mathbf{r}}$ between $M_{p}$ and $M_{p}+1/n$, where $M_{p}$ is the supremum of $\lVert p(X_{1},\dots,X_{d})\rVert$ over all $(X_{1},\dots,X_{d})$ in a tracial $\mathrm{W}^{*}$-algebra $\mathcal{M}$.
Each function and relation symbol is assigned range bounds and moduli of continuity that one would expect, e.g. multiplication is supposed to map $D_{r}\times D_{r^{\prime}}$ into $D_{rr^{\prime}}$ with $\omega_{(D_{r},D_{r}^{\prime}),1}^{\cdot}(t)=r^{\prime}t$ and $\omega_{(D_{r},D_{r}^{\prime}),2}^{\cdot}=rt$. Although not every $\mathcal{L}_{\operatorname{tr}}$-structure comes from a tracial $\mathrm{W}^{*}$-algebra, one can formulate axioms in the language such that any structure satisfying these axioms comes from a tracial $\mathrm{W}^{*}$-algebra [23, §3.2]. For such structures, the notion of homomorphism given above reduces to trace-preserving ***-homomorphisms of tracial $\mathrm{W}^{*}$-algebras.